# Generalizing fusion rules by shuffle: Symmetry-based classifications of nonlocal systems constructed from similarity transformations

### v1

@@ VI Conclusion and discussion
In this paper, we have studied the implication of the Galois shuffle operations in CFTs[83] and their implications as similarity transformations between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems. Essentially, this is a ring isomorphism between two theories, and the RGs of one side are compatible with the other side. However, when studying the implications for defects or boundaries, more careful treatment is necessary, and this difficulty itself is a natural consequence of the nonlocality.
@@ VII Acknowledgement
The authors thank Naomichi Hatano, Hosho Katsura, and Po-Yao Chang for sharing the knowledge on the similarity transformations. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Ingo Runkel and Sylvain Ribault for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on fusion ring. We thank Masaki Oshikawa for the helpful discussions on the large gauge transformation. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315. YF thanks support from NCTS.
@@ Appendix A Generalization and conjecture on similarity transformation
In Hermitian systems, it is known that the Lieb-Shultz-Mattis operators $U_{\alpha}$ which can be realized as an unitary operation changes or shuffle the energy eigenstates. Moreover, they correspond to primary field $\alpha$ and generate many-body excitation from the ground state $U_{\alpha}|0\rangle=\alpha(0)|0\rangle=|\alpha\rangle$ where we have assumed the radial quantization in CFTs and $|0\rangle$ is the vacuum. This unitary operation is called a large gauge transformation and the analogy between the similarity transformation and the gauge transformation has already been pointed out in[120]. Hence, we conjecture the similarity transformation can be implemented as invertible operator $U_{j}$ corresponding to the invertible anyon $j$ or the many-body excitation via the CFT expression $U_{j}|0\rangle=j(0)|0\rangle=|j\rangle$ where $j(0)$ corresponds to a generalized version of LSM operator $U_{j}$ generating the many-body excitations[115, 116, 117, 118].

### v2

@@ VI Conclusion and discussion
In this paper, we have studied the implication of the Galois shuffle operations in CFTs[90] and their implications as similarity transformations between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems. Essentially, this is a ring isomorphism between two theories, and the RGs of one side are compatible with the other side. However, when studying the implications for defects or boundaries, more careful treatment is necessary, and this difficulty is a natural consequence of the nonlocality.
@@ VII Acknowledgement
The authors thank Naomichi Hatano, Hosho Katsura, and Po-Yao Chang for sharing their knowledge on the similarity transformations. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Ingo Runkel and Sylvain Ribault for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on the fusion ring. We thank Masaki Oshikawa for the helpful discussions on the large gauge transformation. YF thanks Shumpei Iino for the discussions on non-GW states. We thank Chen-Te Ma for the helpful comments on the literature on the realization of chiral fermions in non-Hermitian systems. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315. YF tthe hanks support from NCTS.
@@ Appendix A Generalization and conjecture on similarity transformation
In Hermitian systems, it is known that the Lieb-Schultz-Mattis operators $U_{\alpha}$, which can be realized as a unitary operation changes or shuffle the energy eigenstates. Moreover, they correspond tothe primary field $\alpha$ and generate many-body excitation from the ground state $U_{\alpha}|0\rangle=\alpha(0)|0\rangle=|\alpha\rangle$ where we have assumed the radial quantization in CFTs and $|0\rangle$ is the vacuum. This unitary operation is called a large gauge transformation, and the analogy between the similarity transformation and the gauge transformation has already been pointed out in[127]. Hence, we conjecture that the similarity transformation can be implemented as invertible operator $U_{j}$ corresponding to the invertible anyon $j$ or the many-body excitation via the CFT expression $U_{j}|0\rangle=j(0)|0\rangle=|j\rangle$ where $j(0)$ corresponds to a generalized version of LSM operator $U_{j}$ generating the many-body excitations[122, 123, 124, 125].